OpenStudy (anonymous):

Seems simple..but idk why I am getting it wrong :/ I did 3000e^(0.04*1) and I got =3122.43 If you invest $3,000 at 4% interest compounded continuously, what is the average amount in your account over one year? (Round your answer to the nearest cent.)

OpenStudy (anonymous):

what does "average amount" mean? are you supposed to integrate?

OpenStudy (anonymous):

I have no idea that is all it says in the problem, but that would suck if I had to integrate :/

OpenStudy (anonymous):

not really since it is pretty easy to integrate \(e^x\)

OpenStudy (anonymous):

oh then that would just be e^x

OpenStudy (anonymous):

and since the time interval is one year maybe it is \[\int _0^1e^{,04t}dt\]

OpenStudy (anonymous):

times 3000 of course

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

i will try it and see what i get i think that is what it means by "average value' maybe i am wrong

OpenStudy (anonymous):

we put in the 1 and 0 for e?

OpenStudy (anonymous):

or nvm we raise it to those powers right?

OpenStudy (anonymous):

i get 3060.81

OpenStudy (anonymous):

the anti derivative of \(e^{.04t}\) is \(25e^{.04t}\) by a simple u sub

OpenStudy (anonymous):

evaluate at 1 and 0, and subtract

OpenStudy (anonymous):

what did you substitute to get 25e^.04t? sorry I am confused :/

OpenStudy (anonymous):

you get \[3000\times 25\left(e^{.04}-e^0\right)\] or \[75000\left(e^{.04}-1\right)\]

OpenStudy (anonymous):

oh ok you have \(e^{.04t}\) and you want the anti derivative

OpenStudy (anonymous):

the u sub is \(u-.04t, du =.04dt, 25du = dt\) since 25 is the reciprocal of \(.04\)

OpenStudy (anonymous):

i meant \(u=.04t\)

OpenStudy (anonymous):

aka \(u=\frac{1}{25}t\)

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

yes I think I got it!!

OpenStudy (anonymous):

ok good luck if it is an on line class you can check the number i got above and see if it works

OpenStudy (anonymous):

yup it did! I appreciate the help! thanks! :)

OpenStudy (anonymous):

if not, then i am lost oh whew!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)